what is the slope and y-intercept of 3y-x=5
Answer by nerdybill(7384)

what is the slope and y-intercept of 3y-x=5
.
Manipulate the given formula into the "slope-intercept" form of a line:
y = mx + b
where
m is the slope
b is the y-intercept
.
Starting with:
3y-x=5
Add 'x' to both sides:
3y = x+5
Divide both sides by 3:
y = x/3 + 5/3
Or, we can rewrite as:
y = (1/3)x + 5/3
.
So,
slope is 1/3
y-intercept is at (0, 5/3)